OpenStudy (anonymous):

I will fan fan and medal!!!!!!!!!!!!!!!!!! Describe two methods to convert 17/25 to a decimal

OpenStudy (anonymous):

@annas @JackofallTradez

OpenStudy (jackofalltradez):

You could use long division. Or you could use a calculator.

OpenStudy (anonymous):

true...but this is for school...i need a little more description

OpenStudy (jackofalltradez):

Do you know how to describe the long division process? And how you use a calculator to find the decimal?

OpenStudy (anonymous):

not really...i need to review it a little

OpenStudy (here_to_help15):

First, of course, is the division method second is by multiplying the numerator and denominator with a number such that after multiplying, the denominator becomes a multiple of 10 e.g. 17/25 * 4/4 = 68/100 = 0.68 done!

OpenStudy (anonymous):

describe it for 17/25